# Structural dynamics and optimal transport of an active polymer

### v3

@@ Abstract
We study the spontaneous configuration transitions of an active semi-flexible polymer between spiral and non-spiral states, and show that the configuration dynamics is fully described by a subcritical pitchfork bifurcation. Exploiting the fact that active polymer barely moves in spiral states and exhibits net displacements in non-spiral states, we theoretically prove that the motion of the active polymer is consistent with a run-and-tumble-like dynamics. Moreover, we find that there exists an optimal self-propelling force, at which the probabilities of finding the polymer in the spiral and non-spiral state become equal, that maximizes the diffusion coefficient.
@@
Active polymers have recently become an interesting field of research in soft matter and non-equilibrium statistical physics. They play an important role in biological systems such as duplication of DNA by DNA-polymerase albert , synthesizes of proteins by ribosomes albert ; zia , chromatin in eukaryotic cells chromatin1 ; chromatin2 ; chromatin3 ; chromatin4 ; chromatin5 ; chromatin6 , movement of actin filaments by motor proteins in cellular cytoskeletons cytos1 ; cytos2 ; turning6 ; turning7 ; Motility_Assay_Book , microtubule bundles and spools in a kinesin-1 gliding motility assay Chaudhuri_Nano_Lett_2018 , sensing mechanical deformation of the soft materials surface by tracking the gluiding assay of microtubules as active probes on a carpet of kinesin-1 motor proteins Inoue_Nat_Commu_2015 , and the oscillatory beating to drive of flagella flagella0 ; flagella1 ; flagella2 . In all of the above examples, as mechanical tangential forces are generated on the backbone of the objects, they can be considered as active polymers. Motivated by these applications, many studies have focused, both experimentally exp1 ; exp2 ; exp3 ; exp4 ; exp5 and theoretically theo1 ; theo2 ; theo3 ; theo4 ; theo5 ; theo6 ; theo7 ; theo8 ; theo9 ; PRL ; theo10 ; baskaran ; theo11 ; turning ; theo12 ; Jabbari_PRE ; Jabbari_JCP , on active polymers. Theoretical studies have shown that the diffusion coefficient of the flexible/semi-flexible active polymer is a monotonically increasing function – linear or quadratic – of Péclet number theo1 ; theo5 ; theo6 ; theo7 ; theo9 ; PRL ; baskaran ; theo11 ; theo12 .
The structure and configuration of an active polymer have been studied in theo4 ; theo5 ; theo6 ; theo9 ; PRL ; theo10 ; baskaran ; turning ; theo11 ; Jabbari_PRE ; Jabbari_JCP . The interplay between thermal fluctuations and self-propelling stresses (i.e. activity) theo9 ; baskaran ; turning – as occurs when a polymer is embedded in a gliding assay with motor proteins as the source of activity Duke_PRL ; Bourdieu_PRE ; Bourdieu_PRL ; Grill_Nano_Lett ; Tas_Nano_Lett ; Farkas_book ; theo10 ; turning – leads to transient polymer configurations, classified into two main groups: spiral and non-spiral ones. For the latter case turning the radius of gyration (RG) is a convex function with respect to the attachment/detachment rates and Péclet number of the motor proteins, while for a system composed of an active polymer wherein the force is applied to all beads of the polymer theo9 RG is a monotonic decreasing function of the polymer activity [see the Supplemental Material (SM) for more details].
@@
Figure 1: (a) Typical snapshot of a semi-flexible polymer with contour length of $N=200$ and persistence length of $\ell_{\textrm{p}}=5$ after equilibration and just before acting the self-propelling force $F_{\textrm{sp}}$ (orange arrow) on each segment. The polymer is divided to $N/\ell_{\textrm{p}}$ segments. The red and green beads represent the tail and head of each segment, respectively. In each segment $F_{\textrm{sp}}$ acts on the head monomer directing from tail to head (orange dashed line). (b) Definition of the spiral and non-spiral states: probability distribution function of the turning number $P(\psi)$ for fixed values of $F_{\textrm{sp}}=2$, $N=200$ and $\ell_{\textrm{p}}=5$. The vertical black dashed lines at minima of the $P(\psi)$ (located at the left $\psi_{\textrm{min}}^{\textrm{l}}$ and the right $\psi_{\textrm{min}}^{\textrm{r}}$ minima) separate the two spiral and non-spiral states. The polymer with $\psi_{\textrm{min}}^{\textrm{l}}<\psi<\psi_{\textrm{min}}^{\textrm{r}}$ is in the non-spiral state, and it is in the spiral state for $\psi<\psi_{\textrm{min}}^{\textrm{l}}$ or $\psi>\psi_{\textrm{min}}^{\textrm{r}}$. The right and left snapshots correspond to the spiral states with counter-clockwise and clockwise turnings, respectively, and the middle one exhibits a non-spiral state. See SM for movies.
@@
Simulation model – We consider a bead-spring semi-flexible polymer with $N=200$ monomers in a two-dimensional, squared simulation box of area $L\!\times\!L$ – where $L=500\sigma$ and $\sigma$ the length unit – with periodic boundary conditions [Fig. 1(a)]. For non-zero value of the bending rigidity coefficient $\kappa_{\textrm{b}}$, the persistence length of the polymer in two dimensions is obtained as $\ell_{\textrm{p}}=2\kappa_{\textrm{b}}/(k_{\textrm{B}}T)$. The $k_{\textrm{B}}T$ is the thermal energy and $k_{\textrm{B}}$ is the Boltzmann constant [see the SM for more details about the simulation model].
To model the active polymer, the polymer is divided into $N/\ell_{\textrm{{p}}}$ segments. Each segment – see Fig. 1(a) – contains a tail (in red color) and a head (in green color) monomers as the first and last monomers, respectively, and the rest of monomers are in blue color. Employing Langevin dynamics (LD) a SP force $F_{\textrm{sp}}$ acts on the head monomer of each segment and in the direction from tail to head in each time step, depicted by the orange arrow and dashed line, respectively in Fig. 1(a) Amir [see the SM for more details]. All the LD simulations are performed using LAMMPS package lammps .
@@
Figure 2: (a) Non-monotonic diffusion coefficient of the polymer center of mass $D_{\textrm{CM}}$ as a function of the SP force $F_{\textrm{sp}}$, from LD simulations (LD Sim; black circles) and the theory obtained from Eq. (4) (Theory; orange squares). (b) The probability of finding the polymer in the non-spiral state $p_{\textrm{ns}}$ (red squares) and in the spiral state $p_{\textrm{s}}$ (blue circles) as a function of the SP force $F_{\textrm{sp}}$. The intersection of the red and blue curves (denoted by the horizontal black dashed line at $p_{\textrm{ns}}=p_{\textrm{s}}=0.5$ and the vertical black dashed-dotted line at $F_{\textrm{sp}}=1.9$) corresponds to the peak of the black curve $D_{\textrm{CM}}$ in panel (a).
@@
Spiral states & transport – As the snapshots in Fig. 1(b) show, the self-propelled polymer possesses different configurations that correspond to spiral and non-spiral states, which are quantified by considering the turning number $\psi$ of the polymer. The turning number of the entire polymer with $N-1$ bonds is defined as  turn_handbook
@@
where $\theta_{i}$ is the angle between the $i^{\textrm{th}}$ bond vector and horizontal axis. Note that $\psi>0$ [right snapshot in Fig. 1(b)] and $\psi<0$ [left snapshot in Fig. 1(b)] imply that the global configuration of the polymer is counter-clockwise (CCW) and clockwise (CW), respectively. A vanishing $\psi$ value indicates either a rod-like configuration or a random coil polymer configuration [middle snapshot in Fig. 1(b)].
In simulations, we find that the probability distribution function (PDF) of the turning number $P(\psi)$ has three peaks: two are located in the left and right hand sides and one at $\psi=0$; see Fig. 1(b) where two vertical black dashed lines, at the two minima of the $P(\psi)$, indicate the region of $\psi$ that corresponds to non-spiral states. The time evolution $\psi(t)$, the PDF of $P(\psi)$ for various values of $F_{\textrm{sp}}$, and different types of transition between the non-spiral and the spiral states have been explained in the SM. In the spiral state, the sum of all active forces acting on different segments of the polymer is almost negligible and thus the polymer barely moves. On the other hand, when the active polymer escapes from a spiral configuration and stays in the non-spiral state, it is less coiled and active forces along the polymer contribute to displace the center of mass. Fig. 2(a) shows the $D_{\textrm{CM}}$ as a function of $F_{\textrm{sp}}$ for a fixed value of $\ell_{\textrm{p}}=5$ (black circles), where it is evident that $D_{\textrm{CM}}$ is non-monotonic with respect to the $F_{\textrm{sp}}$. The typical trajectories of the polymer CM for various values of $F_{\textrm{sp}}$ and the corresponding mean-square-displacement curves have been presented in the SM. Moreover, it exists an optimal value of $F_{\textrm{sp}}$ that maximizes $D_{\textrm{CM}}$. The maximum of $D_{\textrm{CM}}$ takes place when the probabilities of finding the polymer in the spiral and non-spiral states are equal to each other; see Fig. 2(b).
@@
Figure 3: (a) The PDF of the $\psi$, $P(\psi)$ obtained from integration of Eq. (2) with $\chi=-0.184$ corresponding to SP force $F_{\textrm{sp}}=1.9$ and for fixed value of the noise $\Gamma=0.00375$ (orange circles) and from the LD simulations (purple squared). (b) Bifurcation diagram: theory vs. simulation. The blue solid and dashed lines that are obtained from Eq. (3) show the stable and unstable fixed points, respectively. The filled and empty green circles represent the stable and unstable fixed points, respectively, and are coming from the LD simulations. Inset presents the value of $F_{\textrm{sp}}$ as a function of $\chi$. (c) The PDF of $x$, $P(x)$ (top) and the corresponding potential $\Phi(x)$ (bottom). In both panels the value of $\chi$ has been set to $\chi=-0.2$. (d) The probability of finding the polymer in the non-spiral state $p_{\textrm{ns}}$ (red squares from the LD simulation and red solid line from the theory) and in the spiral state $p_{\textrm{s}}$ (blue circles from the LD simulation and blue solid line from the theory) as a function of $\chi$. The intersection of the red and blue curves (denoted by the horizontal black dashed line at $p_{\textrm{ns}}=p_{\textrm{s}}=0.5$ and the vertical black dashed-dotted line at $\chi=-0.184$ corresponds to $F_{\textrm{sp}}=1.9$) corresponds to the peak of the black curve $D_{\textrm{CM}}$ in Fig. 2(a).
@@
Theory – Next the stability of the turning number is checked by using linear stability analysis strogatz . The PDF of the turning number $\psi$ can be expressed as $P(\psi)\sim\textrm{exp}[-\Phi(\psi)]$ with $\Phi(\psi)=a_{2}\psi^{2}-a_{4}\psi^{4}+a_{6}\psi^{6}$ a potential, where $a_{2},a_{4}$ and $a_{6}$ are constants that depend on the value of $F_{\textrm{sp}}$ and are obtained using the best fitting of the PDF of the turning number coming from LD simulation by the $P(\psi)$ [see Fig. 1(b) and in SM, Fig. 1(b)]; cf. turning . A dynamical system of order $n$th can be quantized by $n$ first order ordinary differential equations strogatz . For the present paper as the turning number $\psi$ is the only dynamical variable of the system, therefore one ordinary differential equation for the $\psi$ in dimensionless units can describe the system as

### v4

@@ Abstract
We study the spontaneous configuration transitions of an active semi-flexible polymer between spiral and non-spiral states, and show that the configuration dynamics is fully described by a subcritical pitchfork bifurcation. Exploiting the fact that an active polymer barely moves in spiral states and exhibits net displacements in non-spiral states, we theoretically prove that the motion of the active polymer is consistent with a run-and-tumble-like dynamics. Moreover, we find that there exists an optimal self-propelling force that maximizes the diffusion coefficient.
I Introduction
@@
Active polymers have recently become an interesting field of research in soft matter and non-equilibrium statistical physics. They play an important role in biological systems such as synthesizes of proteins by ribosomes albert ; zia , chromatin in eukaryotic cells chromatin1 ; chromatin2 ; chromatin3 ; chromatin4 ; chromatin5 ; chromatin6 , movement of actin filaments by motor proteins in cellular cytoskeletons cytos1 ; cytos2 ; turning6 ; turning7 ; Motility_Assay_Book , microtubule bundles and spools in motility assays Chaudhuri_Nano_Lett_2018 , mechanical sensing of soft materials Inoue_Nat_Commu_2015 , directed transport and molecular sorting of microtubules in kinesin-coated nanostructures Dekker_NanoLett_2005 ; Maximov_NanoLett_2013 ; Dekker_Science_2006 , and the coordinated beating of flagella flagella0 ; flagella1 ; flagella2 . In all of the above examples, as mechanical tangential forces are generated on the backbone of the elongated objects, these can be considered as self-propelled polymers. Motivated by these applications, many studies have focused, both experimentally exp1 ; exp2 ; exp3 ; exp4 ; exp5 and theoretically theo1 ; theo2 ; theo3 ; theo4 ; theo5 ; theo6 ; theo7 ; theo8 ; theo9 ; PRL ; theo10 ; baskaran ; theo11 ; turning ; theo12 ; Jabbari_PRE ; Jabbari_JCP , on active polymers. Theoretical studies have shown that the diffusion coefficient of flexible/semi-flexible active polymers is a monotonically increasing function – linear or quadratic – of Péclet number theo1 ; theo5 ; theo6 ; theo7 ; theo9 ; PRL ; baskaran ; theo11 ; theo12 . However, we note that the active forces that are applied on the polymer are not necessary tangential to it. This can occurs, for instance, due to the presence of an active and/or viscoelastic bath ViscoBath or enzymatic force dipoles EnzymActive . While such systems are undoubtedly of great relevance, here our focus is on active polymers driven by tangential forces.
@@
Figure 1: Typical snapshot of a semi-flexible polymer with contour length of $N=200$ and persistence length of $\ell_{\textrm{p}}=5$ after equilibration and just before acting the self-propelling force $F_{\textrm{sp}}$ (orange arrow) on each segment. The polymer is divided to $N/\ell_{\textrm{p}}$ segments. The red and green beads represent the tail and head of each segment, respectively. In each segment $F_{\textrm{sp}}$ acts on the head monomer directing from tail to head (orange dashed line).
Figure 2: (a) The turning number $\psi(t)$ as a function of the time $t$ for fixed values of the polymer contour length of $N=200$, persistence length of $\ell_{\textrm{p}}=5$ and the SP force of $F_{\textrm{sp}}=2$. The top and the bottom insets are typical snapshots of the polymer in the spiral state with $\psi>0$ (CCW) and $\psi<0$ (CW), respectively. In each segment the SP force $F_{\textrm{sp}}$ acts on the green circle directing from the red circle (tail) to the green one (head). (b) Definition of the spiral and non-spiral states: probability distribution function of the turning number $P(\psi)$ for the same values of $F_{\textrm{sp}}$, $N$ and $\ell_{\textrm{p}}$ as of panel (a). The vertical black dashed lines at minima of the $P(\psi)$ (located at the left $\psi_{\textrm{min}}^{\textrm{l}}$ and the right $\psi_{\textrm{min}}^{\textrm{r}}$ minima) separate the two spiral and non-spiral states. The polymer with $\psi_{\textrm{min}}^{\textrm{l}}<\psi<\psi_{\textrm{min}}^{\textrm{r}}$ is in the non-spiral state, and it is in the spiral state for $\psi<\psi_{\textrm{min}}^{\textrm{l}}$ or $\psi>\psi_{\textrm{min}}^{\textrm{r}}$. The right and left snapshots correspond to the spiral states with counter-clockwise and clockwise turnings, respectively, and the middle one exhibits a non-spiral state. See SM for movies.
The structural dynamics of an active polymer has been studied in Refs. 30; 31; 32; 35; 36; 37; 38; 40; 39; 42; 43. The interplay between thermal fluctuations and self-propelling stresses (i.e. activity) theo9 ; baskaran ; turning – as occurs when a polymer is embedded in a gliding assay with motor proteins as the source of activity Duke_PRL ; Bourdieu_PRE ; Bourdieu_PRL ; Grill_Nano_Lett ; Tas_Nano_Lett ; Farkas_book ; theo10 ; turning – leads to transient polymer configurations, classified into two main groups: spiral and non-spiral ones. For the latter case turning the radius of gyration (RG) is a convex function with respect to the attachment/detachment rates and Péclet number of the motor proteins, while for a system composed of an active polymer wherein the force is applied to all beads of the polymer theo9 , the RG is a monotonic decreasing function of the polymer activity [see the Appendix A for more details].
@@
II Simulation model
We consider a bead-spring semi-flexible polymer with $N=200$ monomers in a two-dimensional, squared simulation box of area $L\!\times\!L$ – where $L=500\sigma$ and $\sigma$ the length unit – with periodic boundary conditions (Fig. 1). The $N$ monomers are disks of radius $\sigma$ and mass M. The consecutive monomers are connected by the finitely extensible nonlinear elastic (FENE) potential $U_{\textrm{FENE}}(r)=-\frac{1}{2}kR^{2}_{0}\ln[1-(r/R_{0})^{2}]$, where $r$ is the center-to-center distance between two connected monomers, $k$ the spring constant, and $R_{0}$ the natural length of the spring. Monomers interact via the repulsive Weeks-Chandler-Anderson (WCA) wca potential $U_{\textrm{WCA}}(r)=U_{\textrm{LJ}}(r)-U_{\textrm{LJ}}(r_{\textrm{c}})$ if $r\leq r_{\textrm{c}}$ and zero otherwise, where $r_{c}=2^{1/6}\sigma$ and $r$ are the cut-off radius and the distance between two given monomers, respectively. $U_{\textrm{LJ}}(r)=4\epsilon\big{[}\big{(}\sigma/r\big{)}^{12}-\big{(}\sigma/r\big{)}^{6}\big{]}$ is the Lennard-Jones (LJ) potential, with $\epsilon$ as the depth of the potential well. The rigidity of the semi-flexible polymer is modeled by the cosine potential $U_{\textrm{bend}}(\alpha_{i})=\kappa_{\textrm{b}}[1+\cos\left(\alpha_{i}\right)]$, where $\kappa_{\textrm{b}}$ is a bending constant, $\alpha_{i}$ represents the angle between two consecutive bond vectors connecting $i$th and $i+1$th, and the $i+1$th and $i+2$th beads, whose positive and negative values correspond to the counter-clockwise (CCW) and clockwise (CW) local configurations of the consecutive $i$th and $(i+1)$th bonds, respectively. For a non-zero value of the $\kappa_{\textrm{b}}$, the persistence length of the polymer in two dimensions is obtained as $\ell_{\textrm{p}}=2\kappa_{\textrm{b}}/(k_{\textrm{B}}T)$. The $k_{\textrm{B}}T$ is the thermal energy and $k_{\textrm{B}}$ is the Boltzmann constant. For the sake of (computational) simplicity, the polymer is divided into $N/\ell_{\textrm{{p}}}$ segments. Each segment – see Fig. 1– contains a tail (in red color) and a head (in green color) monomers as the first and last monomers, respectively, and the rest of monomers are in blue color. Importantly, a self-propelling force $F_{\textrm{sp}}$ acts on the head monomer of each segment and in the direction from tail to head, depicted by the orange arrow and dashed line, respectively in Fig. 1 Amir .
Employing Langevin dynamics (LD), the temporal evolution of the $i$th monomer is given by the equation of motion:
$M\ddot{\vec{r}}_{i}=-\gamma\dot{\vec{r}}_{i}+\vec{F}_{\textrm{sp}}\delta_{\textrm{ih}}-\vec{\nabla}_{r}U_{\textrm{tot}}+\sqrt{2\gamma k_{\textrm{B}}T}\vec{\eta}_{i}(t),$ (1)
where $U_{\textrm{tot}}=U_{\textrm{FENE}}+U_{\textrm{bend}}+U_{\textrm{LJ}}$ is the total potential energy and $\gamma$ is the solvent friction coefficient that the $i$th monomer experiences. In the Kronecker delta $\delta_{\textrm{ih}}$ the h stands as an index of the head monomer in each segment. The white noise term follows $\langle\eta_{i}^{m}(t)\rangle=0$ and $\langle\eta_{i}^{m}(t)\eta_{j}^{l}(t^{\prime})\rangle=\delta_{ij}\delta_{ml}\delta(t-t^{\prime})$, where $m,l\equiv x,y$. The M, $\sigma$ and $\epsilon$ are used as the simulation unit scales for mass, length and energy, respectively. The temperature is kept at $k_{\textrm{B}}T=1.2\epsilon$ and the friction coefficient is set to $\gamma=0.7\tau_{0}^{-1}$, in which $\tau_{0}=\sqrt{\textrm{M}\sigma^{2}/\epsilon}$ is the LJ time unit. In our simulations, we consider $dt=0.001\tau_{0}$ and $t_{\textrm{eq}}=5\times 10^{4}\tau_{0}$ as the time step for the integration of the equations of motion and the equilibration time interval, respectively. We also set $k=30$, $R_{0}=1.5$ and $\kappa_{\textrm{b}}=3$ to have a semi-flexible polymer with persistence length of $\ell_{\textrm{p}}=5$. The size of each bead corresponds to the single-strand DNA Kuhn length of $\sigma\approx 1.5$nm. The strength of the interaction at room temperature ($T=295$K) is $3.39\times 10^{-21}$ and the mass of each bead is $936$ amu. Thus, the LJ time scale is 32.1ps. The parameters and variables in the LD simulations are expressed in the LJ units. The polymer is initially equilibrated during the time interval $t_{\textrm{eq}}$ with $F_{\textrm{sp}}=0$. Then the SP force is switched on and the main simulations are done for $10^{6}\tau_{0}$. The results are obtained by averaging over 10 different trajectories, each contains $10^{6}$ data points. All the LD simulations are performed using LAMMPS package lammps .
It is worth stressing that all results presented here hold (qualitatively) by using and over-damped dynamics, instead of Eq. (1): the diffusion coefficient $D_{\textrm{CM}}$ with respect to the SP force $F_{\textrm{sp}}$ exhibits a non-monotonic behavior, the turning number $\psi$ follows the same dynamics, etc.
@@ II Simulation model
Figure 3: (a) Some typical trajectories of the polymer CM for various values of the SP force $F_{\textrm{sp}}=0$ (blue line), 0.5 (red line), 1.9 (green line), 3 (black line) and 10 (orange line). (b) The MSD of the polymer CM as a function of time $t$ for the same values of $F_{\textrm{sp}}$ as of panel (a). The black dashed line $t^{1}$ is a guide to the eye. The non-monotonic behavior in the MSD is denoted by the blue curved arrow. (c) Non-monotonic diffusion coefficient of the polymer center of mass $D_{\textrm{CM}}$ as a function of the SP force $F_{\textrm{sp}}$, from LD simulations (LD Sim; black circles) and the theory obtained from Eq. (5) (Theory; orange squares). (d) The probability of finding the polymer in the non-spiral state $p_{\textrm{ns}}$ (red squares) and in the spiral state $p_{\textrm{s}}$ (blue circles) as a function of the SP force $F_{\textrm{sp}}$. The intersection of the red and blue curves (denoted by the horizontal black dashed line at $p_{\textrm{ns}}=p_{\textrm{s}}=0.5$ and the vertical black dashed-dotted line at $F_{\textrm{sp}}=1.9$) corresponds to the peak of the black curve $D_{\textrm{CM}}$ in panel (c).
III Spiral states & transport
@@
The self-propelled polymer possesses different configurations that correspond to spiral and non-spiral states, which are quantified by considering the turning number $\psi$ of the polymer. The turning number of the entire polymer with $N-1$ bonds is defined as  turn_handbook
@@
where $\theta_{i}$ is the angle between the $i^{\textrm{th}}$ bond vector and horizontal axis. In Fig. 2(a) the time evolution of the turning number $\psi(t)$ has been plotted. The top left, top right and the bottom insets of Fig. 2 are typical snapshots of the polymer in the non-spiral state, spiral state with $\psi>0$ (CCW) and $\psi<0$ (CW), respectively. As mentioned before, in each segment the SP force $F_{\textrm{sp}}$ acts on the green circle directing from the red circle (tail) to the green one (head). Note that $\psi>0$ [top right snapshot in Fig. 2(b)] and $\psi<0$ [bottom left snapshot in Fig. 2(a)] imply that the global configuration of the polymer is counter-clockwise (CCW) and clockwise (CW), respectively. A vanishing $\psi$ value indicates either a rod-like configuration or a random coil polymer configuration [top left snapshot in Fig. 2(a)].
In simulations, we find that the probability distribution function (PDF) of the turning number $P(\psi)$ has three peaks: two are located in the left and right hand sides and one at $\psi=0$; see Fig. 2(b) where two vertical black dashed lines, at the two minima of the $P(\psi)$, indicate the region of $\psi$ that corresponds to non-spiral states. The dependency of $P(\psi)$ with the active force $F_{\textrm{sp}}$ is shown in Appendix B. The different types of transition between the non-spiral and the spiral states are displayed and discussed in Appendix C.
In the spiral state, the sum of all active forces acting on different segments of the polymer is almost negligible and thus the center of mass of the polymer barely moves compared to the non-spiral state. On the other hand, when the active polymer escapes from a spiral configuration (and stays in the non-spiral state), it is less coiled and active forces along the polymer contribute significantly more to the displacement of the center of mass. Fig. 3(a) illustrates some typical trajectories of the polymer CM for different values of the SP force $F_{\textrm{sp}}=0$ (in blue color), 0.5 (in red color), 1.9 (in green color), 3 (in black color) and 10 (in orange color). Note that the the covered area by the polymer center of mass in a given time first increases and then decreases with the SP force. This is confirmed by the non-monotonic behavior in the MSD as a function of time; see Fig. 3(b). The superdiffusive regime due to the polymer activity at the intermediate time scales is followed by the diffusive regime at long time scales. For large values of the $F_{\textrm{sp}}$ an oscillatory behavior in the MSD at the intermediate time scales is seen (orange line for $F_{\textrm{sp}}=10$) at which the polymer spends most of its time in the spiral state.
Fig. 3(c) shows the $D_{\textrm{CM}}$ as a function of $F_{\textrm{sp}}$, computed from the mean squared displacement (MSD) curves in panel (b), for a fixed value of $\ell_{\textrm{p}}=5$ (black circles), where it is evident that $D_{\textrm{CM}}$ is non-monotonic with respect to the $F_{\textrm{sp}}$. We find that there exists an optimal value of $F_{\textrm{sp}}$ that maximizes $D_{\textrm{CM}}$. The maximum of $D_{\textrm{CM}}$ takes place when the probabilities of finding the polymer in the spiral state $P_{\textrm{s}}$ and non-spiral state $P_{\textrm{ns}}$ are equal to each other; see Fig. 3(d).
@@
Figure 4: (a) The PDF of the $\psi$, $P(\psi)$ obtained from integration of Eq. (3) with $\chi=-0.184$ corresponding to SP force $F_{\textrm{sp}}=1.9$ and for fixed value of the noise $\Gamma=0.00375$ (orange circles) and from the LD simulations (purple squared). (b) Bifurcation diagram: theory vs. simulation. The blue solid and dashed lines that are obtained from Eq. (4) show the stable and unstable fixed points, respectively. The filled and empty green circles represent the stable and unstable fixed points, respectively, and are coming from the LD simulations. Violet stars denote two additional saddle-node bifurcations at $\chi_{\textrm{s}}=-1/4$. Inset presents the value of $F_{\textrm{sp}}$ as a function of $\chi$. (c) The PDF of $x$, $P(x)$ (top) and the corresponding potential $\Phi(x)$ (bottom). In both panels the value of $\chi$ has been set to $\chi=-0.2$. (d) The probability of finding the polymer in the non-spiral state $p_{\textrm{ns}}$ (red squares from the LD simulation and red solid line from the theory) and in the spiral state $p_{\textrm{s}}$ (blue circles from the LD simulation and blue solid line from the theory) as a function of $\chi$. The intersection of the red and blue curves (denoted by the horizontal black dashed line at $p_{\textrm{ns}}=p_{\textrm{s}}=0.5$ and the vertical black dashed-dotted line at $\chi=-0.184$ corresponds to $F_{\textrm{sp}}=1.9$) corresponds to the peak of the black curve $D_{\textrm{CM}}$ in Fig. 3(c).
@@
IV Theory
To analyze the structural dynamics of the polymer, we observe that in the steady state the PDF of the turning number $\psi$, $P(\psi)$, can be empirically described by $P(\psi)\sim\textrm{exp}[-\Phi(\psi)]$ with $\Phi(\psi)=a_{2}\psi^{2}-a_{4}\psi^{4}+a_{6}\psi^{6}$ a potential, where $a_{2},a_{4}$ and $a_{6}$ are constants that depend on the value of $F_{\textrm{sp}}$ and are obtained using the best fitting of the PDF of the turning number coming from LD simulation by the $P(\psi)$ [see Fig. 2(b), and also Fig. 6 in Appendix B]; cf. Ref. 40. Let us recall that given a Langevin equation $\dot{x}=-\frac{\partial u}{\partial x}+\xi$, with $\xi$ a standard white noise (with intensity $1$), the steady state solution of the corresponding Fokker-Planck equation is $P(x)\sim\textrm{exp}[-u(x)]$. We use this results to expressed the dynamics of $\psi$, in dimensionless units, as: